# Cross-Modal Causal Intervention for Medical Report Generation

### v4

@@ III-B Visual-Linguistic Pre-training (VLP)
In the medical pre-training framework, there exist two difficulties: (1) the unpaired data that only has a single modality is hard to be utilized in supervised learning, and (2) heterogeneous data that makes it difficult to distinguish the region features because the morphology of the same lesion varies greatly [2]. To address these challenges and provide fine-grained region features without region labels, we employ a cross-modal pre-training approach to learn and align multi-modal information [47], as shown in Fig. 2(c). For language concepts, we employ Prefix Language Modeling (PLM) to align with the visual modality utilizing bidirectional attention in image and prefix text. Meanwhile, to acquire visual context, we utilize Masked Image Modeling (MIM) for pixel-level semantic learning. Additionally, to capture more detailed features, we utilize degradation of the image and masking the features extracted by CNN. Therefore, we use a multiway transformer to extract multi-modal features and two linear layers to solve PLM and MIM tasks, respectively [28, 48]. In each block of the multiway encoder, the attention layer is weight-shared while the two feed-forward layers handle the corresponding modal features respectively [49]. Similarly, each block of the multiway decoder consists of a weight-shared self-attention layer, a pool of feed-forward networks used for different modalities, as shown in Fig. 2(b).
III-B1 PLM
@@
Motivated by the work of SimVLM [28], we extract image features from the first three blocks of ResNet101 [50] as prefix tokens in the PLM. Simultaneously, the text is randomly divided into two parts, with one part generated by another under the guidance of the obtained image tokens. In cases where the corresponding image is missing, the PLM can still be trained using only the text, similar to SimVLM. Let denote the image token extracted from the raw image , where represents the patch size and is the embedding size. Then, represents the postfix sequence following the textual description of length . Thus, the formulation is as follows:
@@
III-B2 MIM
To handle unpaired images, we leverage the Masked Image Modeling (MIM) paradigm [48]. Furthermore, since MIM can be trained using pairwise data [51], the missing semantics of masked images can be complemented by text, thereby enhancing cross-modal association. Additionally, we learn medical image representations by reconstructing high-resolution patches from low-resolution inputs, which can encode more local information into latent embeddings [52]. Consequently, we sub-sample the images for degradation before the visual embedding and reconstruct the masked visual token extracted from CNNs by incorporating the semantics of both the unmasked visual token and the linguistic token. This approach enables us to capture subtle differences in the dataset [53]. The objective of MIM can be formulated as:
@@ III-C Visual-Linguistic Causal Intervention
After the Visual-Linguistic pre-training (VLP), the learned visual-linguistic feature encoders still contain visual and linguistic biases from cross-modal confounders [17]. Therefore, we further employ front-door causal intervention to discover the causal effect between visual and linguistic modalities during MRG, as shown in Fig. 2(d).
@@ III-C1 Preliminaries
where the confounders generally brings about the observational bias via [14]. To alleviate this issue, back-door causal intervention can be implemented by introducing the do calculus [14, 15, 54, 55]. We can calculate the probabilities of observable confounders and block the back-door path , the interventional probability is as follows:
where can be the learned RoI features of the heart and the attended word feature of “Normal Heart”. However, the back-door causal intervention is limited by the observability of confounders, and the front-door causal intervention provides an implicit method of deconfounding in Fig. 3 (d). To eliminate the unobservable confounder, we introduce the mediator to cut off the link . The total probability can be represented as the following summation:
@@
However, existing front-door causal intervention methods employ diverse approaches for estimating mediators and conducting interventions. In MRG, we construct the structural causal model (SCM) similar to Image Caption [14], assuming that is the visual feature, is the linguistic feature of the attended word, and is the fusion feature from and (), leading to the generation of report . Moreover, would be influenced by through cross-attention and makes , as shown in Fig. 3 (e). Following that, we divide the mediator into visual and linguistic modalities and introduce a Front-door Intervention Module (FIM) based on Equation (8) to deconfound them.
@@ III-C2 Front-door Intervention Module (FIM)
where brings the spurious correlation via , leading to incorrect reports. is the visual token from visual embedding, and is the linguistic token of the attended word from linguistic embedding. Taking Fig. 1 as an example, when is large while is small in the training set, it tends to enlarge in the testing set. To mitigate visual-linguistic confounders and uncover the true causal structure, we apply causal front-door intervention by introducing mediator and , respectively, as shown in Fig. 3 (f). Generally, is unobservable without a well-trained object detector, and the back-door path can be blocked by via learning the true causal effect . Similarly, the intervention on the back-door path can be implemented by calculating the without well-constructed confounders dictionaries. Thus, we can formulate Eq. (9) as:
@@
where denote the network mapping functions, and denote the estimations of and via VDM and LDM. In Fig. 4, the FIM consists of two Attention Fusion layers and it is integrated into VDM and LDM. Different from the cascaded transformer [25], FIM does not have optimized parameters. It relies on extracted features and intervenes using estimated mediators of the corresponding modalities. This approach ensures that the estimation of mediators is entirely dependent on the sampling in VDM and LDM rather than FIM, achieving implicit learning of mediators.
@@ III-C3 Visual Deconfounding Module (VDM)
Local Sampling. Inspired by TransFG [58], we leverage the attention accumulated from the encoder to select the top tokens. These selected visual tokens with high attention correspond to the report’s keywords as , where for each attention head, and is the dimension of the transformer. Subsequently, is further enhanced using CaaM [32], which excavates the local internal relations. Specifically, in CaaM, these highly attended tokens are computed not only with self-attention but also with negative attention scores, aiming to achieve more robust features. The purpose of is to capture crucial local details in the image, which serve as the key basis for further processing.
@@ IV-A Experimental settings
| R2Gen | 97.14 | 78.07 | 3.66 | 0.103 |
| R2GenCMN | 40.25 | 58.65 | 12.98 | 0.105 |
| VLCI | 97.10 | 69.41 | 10.50 | 0.119 |
IV-A1 Dataset
IU-Xray [63], also known as the Indiana University Chest X-ray Collection, is a publicly available medical dataset widely used to evaluate the performance of MRG methods. It comprises 7,470 chest images and 3,955 corresponding reports, with an average report length of 30 [39]. To maintain consistency, we follow the setting used in R2Gen [5] and partition the dataset into training, validation, and testing sets at a ratio of 7:1:2, ensuring that there is no overlap in patients. We tokenize the words with more than 3 occurrences and set the max length as 60. Note that we adopt two images (frontal and lateral views) as input for each sample.
MIMIC-CXR [64] is a large-scale chest medical dataset, with 377,110 images and 227,835 corresponding reports, of which the average length of reports is 48 in the training/val set and 61 in testing set [39]. We use the official paradigm, the dataset is divided into the training set with 368,960 images and 222,758 reports, the validation set with 2,991 images and 1,808 reports, and the testing set with 5,159 images and 3,269 reports. Different from the IU-Xray dataset, MIMIC-CXR has instances of a single modality (only image or report) as well as multiple images corresponding to one report. Therefore, we use a single image as the input and tokenize the words with more than 10 occurrences, and set the max length as 80.
| Method | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | CIDEr | Rouge-L | METEOR | Precision | Recall | F1 |
| IU-Xray Dataset | ||||||||||
| R2Gen[5] | 0.470 | 0.304 | 0.219 | 0.165 | - | 0.371 | 0.187 | - | - | - |
| CMCL[44] | 0.473 | 0.305 | 0.217 | 0.162 | - | 0.378 | 0.186 | - | - | - |
| PPKED[7] | 0.483 | 0.315 | 0.224 | 0.168 | 0.351 | 0.376 | 0.190 | - | - | - |
| CA[8] | 0.492 | 0.314 | 0.222 | 0.169 | - | 0.381 | 0.193 | - | - | - |
| AlignTransformer[42] | 0.484 | 0.313 | 0.225 | 0.173 | - | 0.379 | 0.204 | - | - | - |
| M2TR[45] | 0.486 | 0.317 | 0.232 | 0.173 | - | 0.390 | 0.192 | - | - | - |
| MGSK[9] | 0.496 | 0.327 | 0.238 | 0.178 | 0.382 | 0.381 | - | - | - | - |
| RAMT[62] | 0.482 | 0.310 | 0.221 | 0.165 | - | 0.377 | 0.195 | - | - | - |
| MMTN[61] | 0.486 | 0.321 | 0.232 | 0.175 | 0.361 | 0.375 | - | - | - | - |
| DCL[40] | - | - | - | 0.163 | 0.586 | 0.383 | 0.193 | - | - | - |
| 0.505 | 0.334 | 0.245 | 0.190 | 0.592 | 0.394 | 0.210 | - | - | - | |
| MIMIC-CXR Dataset | ||||||||||
| R2Gen[5] | 0.353 | 0.218 | 0.145 | 0.103 | 0.149 | 0.277 | 0.142 | 0.333 | 0.273 | 0.276 |
| R2GenCMN[46] | 0.356 | 0.219 | 0.147 | 0.105 | 0.152 | 0.278 | 0.141 | 0.334 | 0.275 | 0.278 |
| CMCL[44] | 0.334 | 0.217 | 0.140 | 0.097 | - | 0.281 | 0.133 | - | - | - |
| PPKED[7] | 0.360 | 0.224 | 0.149 | 0.106 | 0.237 | 0.284 | 0.149 | - | - | - |
| CA[8] | 0.350 | 0.219 | 0.152 | 0.109 | - | 0.283 | 0.151 | 0.352 | 0.298 | 0.303 |
| AlignTransformer[42] | 0.378 | 0.235 | 0.156 | 0.112 | - | 0.283 | 0.158 | - | - | - |
| M2TR[45] | 0.378 | 0.232 | 0.154 | 0.107 | - | 0.272 | 0.145 | 0.240 | 0.428 | 0.308 |
| MGSK[9] | 0.363 | 0.228 | 0.156 | 0.115 | 0.203 | 0.284 | - | 0.458 | 0.348 | 0.371 |
| RAMT[62] | 0.362 | 0.229 | 0.157 | 0.113 | - | 0.284 | 0.153 | 0.380 | 0.342 | 0.335 |
| MMTN[61] | 0.379 | 0.238 | 0.159 | 0.116 | - | 0.283 | 0.161 | - | - | - |
| DCL[40] | - | - | - | 0.109 | 0.281 | 0.284 | 0.150 | 0.471 | 0.352 | 0.373 |
| 0.400 | 0.245 | 0.165 | 0.119 | 0.190 | 0.280 | 0.150 | 0.489 | 0.340 | 0.401 | |
@@ IV-A2 Baseline Models
We compare the proposed VLCI model with several state-of-the-art MRG models, including R2Gen [5], R2GenCMN [46], CMCL [44], PPKED [7], CA [8], AlignTransformer [42], M2TR [45], RAMT [62], MMTN [61], MGSK [9], and DCL [40]. These models are categorized into lightweight and heavyweight models. The lightweight models comprise no more than 3-layer modules for both the encoder and the decoder, as detailed in Table I. Most of these models incorporate various modules to enhance their performance, such as the knowledge-aware module, template retrieval module, and memory-drive module, which can be computationally expensive. It is important to note that the total parameters of our VLCI are less than the R2Gen, while VLCI is more efficient by eliminating the recursive memory calculation dependency, as shown in Table II.
@@ IV-A3 Evaluation Metrics
We adopt the widely used natural language generation (NLG) metrics, including BLEU [65], ROUGE-L [66], METEOR [67] and CIDEr [68]. Since the MRG specifically focuses on the abnormality detection accuracy rather than the text fluency and similarity with the real report, we further adopt clinical efficacy (CE) metrics [5, 8, 45, 9]. It is calculated by the labels extracted from CheXpert [69]. Specifically, the extracted positive labels are considered as positives, while the non-positive labels (negative, not mentioned, and uncertain) are treated as negatives. Using this approach, we calculate the micro-precision, micro-recall, and micro-F1 scores between the labels from reference reports and generated reports.

### v5

@@ III-B RadCARE
In the radiological pre-training framework, there exist two difficulties: (1) the unpaired data that only has a single modality is hard to be utilized in supervised learning, and (2) heterogeneous data that makes it difficult to distinguish the region features because the morphology of the same lesion varies greatly [2]. To address these challenges and provide fine-grained region features without region labels, we employ a visual-linguistic pre-training approach to learn and align visual-linguistic information, as shown in Fig. 2 (b). Unlike existing methods that simply combine Masked Language Modeling (MLM) with Masked Image Modeling (MIM) [60] or rely on contrastive learning for cross-modal alignment [61, 62, 63], we integrate postfix text generation task with degradation-aware masked image restoration task to achieve efficient cross-modal representation learning and alignment. To address the issue of missing modalities, we implement alternating training using multiple data input schemes, surpassing existing approaches that rely on a single training strategy. Moreover, to effectively capture the subtle details in radiological images, we employ a degradation-aware mechanism for pixel-level cross-modal semantic learning, thereby enhancing the alignment of visual and linguistic information.
Specifically, RadCARE introduces four strategies for pre-training as shown in Fig 3 (a), including 1) generation of postfix text with text prefixes as input, 2) restoration of the masked image with both complete text and degraded images as input, 3) generation of postfix text with degraded images and text prefixes as input, and 4) restoration of the masked image with degraded images as input. To enable flexible switching between various strategies, we use a multiway transformer to extract multi-modal features and two linear layers to solve text generation and image restoration tasks, respectively [18, 17], as shown in Fig2 (b). In each block of the multiway encoder, the attention layer is weight-shared while the two feed-forward layers handle the corresponding modal features respectively [64]. Similarly, each block of the multiway decoder consists of a weight-shared self-attention layer, a pool of feed-forward networks used for different modalities, as shown in Fig. 3 (b). Additionally, the noise present in the MIMIC-CXR dataset, such as tasks requiring differentiation between the left and right lung despite being given lateral views, hinders the model’s ability to learn robustly [65, 66]. To address this issue, we combined two datasets for pre-training, as the IU-Xray dataset contains both lateral and frontal views.
III-B1 Postfix Text Generation
@@
Motivated by the work of SimVLM [18], we extract image features from the first three blocks of ResNet101 [67] as prefix tokens for the RadCARE. Simultaneously, the text is randomly divided into two parts, with one part generated by another under the guidance of the obtained image tokens. In cases where the corresponding image is missing, the RadCARE can still be trained using only the text. Let denote the image token extracted from the raw image , where represents the patch size and is the embedding size. Then, represents the postfix sequence following the textual description of length . Thus, the formulation is as follows:
@@
III-B2 Masked Image Restoration
To handle unpaired images, we leverage the MIM paradigm [17]. Furthermore, since masked image restoration can be trained using pairwise data [68], the missing semantics of masked images can be complemented by text, thereby enhancing cross-modal association. Additionally, we learn radiological image representations by reconstructing high-resolution patches from low-resolution inputs, which can encode more local information into latent embeddings [69]. Consequently, we sub-sample the images for degradation before the visual embedding and reconstruct the masked visual token extracted from CNNs by incorporating the semantics of both the unmasked visual token and the linguistic token. This degradation-aware approach enables us to capture subtle differences in the dataset [70, 71]. The objective of masked image restoration can be formulated as:
@@ III-C Visual-Linguistic Causal Intervention
After the RadCARE, the learned visual-linguistic feature encoders still contain visual and linguistic biases from cross-modal confounders [15]. Therefore, we further employ visual-linguistic causal intervention (VLCI) to discover the causal effect between visual and linguistic modalities during RRG, as shown in Fig. 2 (b).
@@ III-C1 Preliminaries
where the confounders generally brings about the observational bias via [26]. To alleviate this issue, back-door causal intervention can be implemented by introducing the do calculus [26, 13, 72, 73]. In Fig. 4 (c), we can calculate the probabilities of observable confounders and block the back-door path , the interventional probability is as follows:
where can be the learned RoI features of the heart and the attended word feature of “Normal Heart”. Based on the conditional probabilities following intervention, it can estimate the true causal effects of each factor in the absence of confounding influences. These individual causal effects can then be aggregated to derive the overall causal effect under the given input conditions. However, the back-door causal intervention is limited by the observability of confounders, and the front-door causal intervention provides a method of deconfounding in Fig. 4 (d). To eliminate the unobservable confounder, we introduce the mediator to cut off the link . The total probability can be represented as the following summation:
@@
However, existing front-door causal intervention methods employ diverse approaches for estimating mediators and conducting interventions. In RRG, we construct the structural causal model (SCM) similar to Image Caption [26], assuming that is the visual feature, is the linguistic feature of the attended word, and is the fusion feature from and (), leading to the generation of report . Moreover, would be influenced by through cross-attention and makes , as shown in Fig. 4 (e).
Previous method [26] employed a pre-constructed semantic dictionary for backdoor intervention, without accounting for complex representations found in radiological imaging. Differently, we construct mediators internally within the model to simulate front-door causal interventions, as illustrated in Figure 7. We further mine and construct mediators separately from visual and linguistic aspects by leveraging semantic correlations obtained during the pre-training phase. Finally, we estimate the actual causal effects by conditioning each mediator through mediator-based sub-distributions. In contrast, the baseline merely relies on the confounded distribution for estimation. To address this, we introduce the Front-Door Intervention Module (FIM) based on Equation (8) to disentangle the effects and construct the Visual Mediator Module (VDM) and Linguistic Mediator Module (LDM) separately.
@@ III-C2 Front-door Intervention Module (FIM)
where brings the spurious correlation via , leading to incorrect reports. is the visual token from visual embedding, and is the linguistic token of the attended word from linguistic embedding.
Taking Fig. 1 as an example, when is large while is small in the training set, it tends to enlarge in the testing set.
To mitigate visual-linguistic confounders and uncover the true causal structure, we apply causal front-door intervention by introducing mediator and , respectively, as shown in Fig. 4 (f). Generally, is unobservable without a well-trained object detector, and the back-door path can be blocked by via learning the true causal effect . Similarly, the intervention on the back-door path can be implemented by calculating the without well-constructed confounders dictionaries. Thus, we can formulate Eq. (9) as:
@@
where denote the network mapping functions including the transformer decoder and a linear layer, and denote the estimations of and via VDM and LDM. In Fig. 5, the FIM consists of two Attention Fusion layers and it is integrated into VDM and LDM. Different from the cascaded transformer [25], FIM does not have optimized parameters. It relies on extracted features and intervenes using estimated mediators of the corresponding modalities. This approach ensures that the estimation of mediators is entirely dependent on the sampling in VDM and LDM rather than FIM, achieving learning of mediators.
@@ III-C3 Visual Deconfounding Module (VDM)
Local Sampling. We leverage the attention accumulated from the encoder to select the top tokens [76]. These selected visual tokens with high attention correspond to the report’s keywords as , where for each attention head, and is the dimension of the transformer. Subsequently, is further enhanced using CaaM [37], which excavates the local internal relations. Specifically, these highly attended tokens are computed not only with self-attention but also with negative attention scores, aiming to achieve more robust features. The purpose of is to capture crucial local details in the image, which serve as the key basis for further processing.
@@ IV-A Experimental settings
IV-A1 Datasets
IU-Xray [81], also known as the Indiana University Chest X-ray Collection, is a publicly available radiological dataset widely used to evaluate the performance of RRG methods. It comprises 7,470 chest images and 3,955 corresponding reports, with an average report length of 30. To maintain consistency, we follow the setting used in R2Gen [5] and partition the dataset into training, validation, and testing sets at a ratio of 7:1:2, ensuring that there is no overlap in patients. We tokenize the words with more than 3 occurrences and set the max length as 60. Note that we adopt two images (frontal and lateral views) as input for each sample.
MIMIC-CXR [82] is a large-scale chest radiological dataset, with 377,110 images and 227,835 corresponding reports, of which the average length of reports is 48 in the training/val set and 61 in testing set. We use the official paradigm, the dataset is divided into the training set with 368,960 images and 222,758 reports, the validation set with 2,991 images and 1,808 reports, and the testing set with 5,159 images and 3,269 reports. Different from the IU-Xray dataset, MIMIC-CXR has instances of a single modality (only image or report) as well as multiple images corresponding to one report. Therefore, we use a single image as the input and tokenize the words with more than 10 occurrences, and set the max length as 80.
@@ IV-A1 Datasets
R2Gen | 97.14 | 78.07 | 3.66 | 10.3 |
R2GenCMN | 40.25 | 58.65 | 12.98 | 10.5 |
CMCRL | 97.10 | 69.41 | 10.50 | 11.9 |
Method | BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | Rouge-L | METEOR | Precision | Recall | F1 |
IU-Xray Dataset | |||||||||
R2Gen[5] | 47.0 | 30.4 | 21.9 | 16.5 | 37.1 | 18.7 | - | - | - |
CMCL[52] | 47.3 | 30.5 | 21.7 | 16.2 | 37.8 | 18.6 | - | - | - |
PPKED[8] | 48.3 | 31.5 | 22.4 | 16.8 | 37.6 | 19.0 | - | - | - |
CA[9] | 49.2 | 31.4 | 22.2 | 16.9 | 38.1 | 19.3 | - | - | - |
AlignTransformer[50] | 48.4 | 31.3 | 22.5 | 17.3 | 37.9 | 20.4 | - | - | - |
M2TR[54] | 48.6 | 31.7 | 23.2 | 17.3 | 39.0 | 19.2 | - | - | - |
MGSK[10] | 49.6 | 32.7 | 23.8 | 17.8 | 38.1 | - | - | - | - |
RAMT[80] | 48.2 | 31.0 | 22.1 | 16.5 | 37.7 | 19.5 | - | - | - |
MMTN[79] | 48.6 | 32.1 | 23.2 | 17.5 | 37.5 | - | - | - | - |
DCL[49] | - | - | - | 16.3 | 38.3 | 19.3 | - | - | - |
SSVE[55] | 49.2 | 32.1 | 23.3 | 18.0 | 37.9 | 20.3 | - | - | - |
Med-LLM[57] | - | - | - | 16.8 | 38.1 | 20.9 | - | - | - |
50.5 | 33.4 | 24.5 | 19.0 | 39.4 | 21.0 | - | - | - | |
MIMIC-CXR Dataset | |||||||||
R2Gen[5] | 35.3 | 21.8 | 14.5 | 10.3 | 27.7 | 14.2 | 33.3 | 27.3 | 27.6 |
R2GenCMN[56] | 35.6 | 21.9 | 14.7 | 10.5 | 27.8 | 14.1 | 33.4 | 27.5 | 27.8 |
CMCL[52] | 33.4 | 21.7 | 14.0 | 09.7 | 28.1 | 13.3 | - | - | - |
PPKED[8] | 36.0 | 22.4 | 14.9 | 10.6 | 28.4 | 14.9 | - | - | - |
CA[9] | 35.0 | 21.9 | 15.2 | 10.9 | 28.3 | 15.1 | 35.2 | 29.8 | 30.3 |
AlignTransformer[50] | 37.8 | 23.5 | 15.6 | 11.2 | 28.3 | 15.8 | - | - | - |
M2TR[54] | 37.8 | 23.2 | 15.4 | 10.7 | 27.2 | 14.5 | 24.0 | 42.8 | 30.8 |
MGSK[10] | 36.3 | 22.8 | 15.6 | 11.5 | 28.4 | - | 45.8 | 34.8 | 37.1 |
RAMT[80] | 36.2 | 22.9 | 15.7 | 11.3 | 28.4 | 15.3 | 38.0 | 34.2 | 33.5 |
MMTN[79] | 37.9 | 23.8 | 15.9 | 11.6 | 28.3 | 16.1 | - | - | - |
DCL[49] | - | - | - | 10.9 | 28.4 | 15.0 | 47.1 | 35.2 | 37.3 |
CAMANet[53] | 37.4 | 23.0 | 15.5 | 11.2 | 27.9 | 14.5 | 48.3 | 32.3 | 38.7 |
40.0 | 24.5 | 16.5 | 11.9 | 28.0 | 15.0 | 48.9 | 34.0 | 40.1 |
@@ IV-A2 Baseline Models
We compare the proposed CMCRL model with several state-of-the-art RRG models, including R2Gen [5], R2GenCMN [56], CMCL [52], PPKED [8], CA [9], AlignTransformer [50], M2TR [54], RAMT [80], MMTN [79], MGSK [10], DCL [49], SSVE[55], CAMANet[53], and Med-LMM[57]. These models are categorized into lightweight and heavyweight models. The lightweight models comprise no more than 3-layer modules for both the encoder and the decoder, as detailed in Table I. Most of these models incorporate various modules to enhance their performance, such as the knowledge-aware module, template retrieval module, and memory-drive module, which can be computationally expensive. It is important to note that the total parameters of our CMCRL are less than the R2Gen, while CMCRL is more efficient by eliminating the recursive memory calculation dependency, as shown in Table II.
@@ IV-A3 Evaluation Metrics
We adopt the widely used natural language generation (NLG) metrics, including BLEU [83], a metric evaluates how similar the generated text is to reference texts by calculating n-gram precision. ROUGE-L [84] evaluates both precision and recall and also considers synonyms, stemming, and paraphrasing. METEOR [85] focuses on the longest common subsequence between the generated text and the reference text, capturing fluency and coherence. CIDEr [86] considers term frequency-inverse document frequency (TF-IDF) weighting to reduce the impact of overly common words. Since the RRG specifically focuses on the abnormality detection accuracy rather than the text fluency and similarity with the real report, we further adopt clinical efficacy (CE) metrics [5, 9, 54, 10]. It is calculated by the labels extracted from CheXpert [87]. Specifically, the extracted positive labels are considered as positives, while the non-positive labels (negative, not mentioned, and uncertain) are treated as negatives. Using this approach, we calculate the micro-precision, micro-recall, and micro-F1 scores between the labels from reference reports and generated reports.